Lulabalula:

Salina spent 100 minutes finishing her homework last night. She spent 23% of her time completing her math homework. Which model represents 23%? A grid model with 100 squares. 13 squares are shaded. A grid model with 100 squares. 23 squares are shaded. A grid model with 100 squares. 27 squares are shaded. A grid model with 100 squares. 73 squares are shaded.

Convert:

23%= 23/100

Lulabalula:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Convert 23%= 23/100 \(\color{#0cbb34}{\text{End of Quote}}\) yeah

Convert:

so which option would show 23% shaded if 23% = 23/100

Lulabalula:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Convert so which option would show 23% shaded if 23% = 23/100 \(\color{#0cbb34}{\text{End of Quote}}\) the second one

Convert:

Yes that's what I would think it'd be